OpenStudy (anonymous):

what are all the real zeroes y=(x-12)^3-7

OpenStudy (amistre64):

it would be nice if we could somehow get rid of that -12 and insert a cuberoot(7)

OpenStudy (anonymous):

how do you do that

OpenStudy (anonymous):

This is what I have but not sure if right y=(x-12)-3 sgrt 7(x-12)^2+(x-12)3sgrt7+3sgrt49 x=12+3sgrt7

OpenStudy (amistre64):

that is correct, yes; just make sure that 3sgrt is meant to be a cuberoot radical